OpenStudy (anonymous):

Frank jogs in a circular garden every day. The ground measures 5 feet from the center to the inner circle and an extra 3 feet from the inner circle to the outer circle. If Frank jogs around the outer circle twice, how much distance does he cover?

OpenStudy (anonymous):

@elevel17

OpenStudy (anonymous):

@KenLJW

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):



OpenStudy (primeralph):

tell them to get a treadmill. Quite cheap nowadays.

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

The radius of the outer circle is 8 ft. You need the circumference of the circle. C = 2(pi)r Plug in 8 ft for r. Then multiply the result by two since Frank goes twice around the circle.

OpenStudy (anonymous):

hmmm well i got 50.24 but i think its asking for something else couse the answers possible are A. 2p(5 + 8) B. 2p(10 + 16) C. p(8 – 5) D. 2p(8 + 8)

OpenStudy (anonymous):

O! so its A!

OpenStudy (anonymous):

Thanks for the help been working on this problem forever

OpenStudy (mathstudent55):

It's D

OpenStudy (anonymous):

how?

OpenStudy (mathstudent55):

C = 2(pi)r = 2 *pi * 8 for one lap Two laps are 2pi8 + 2pi8 = 2pi(8 + 8)

OpenStudy (whpalmer4):

\(2\pi(5+8)\) represents the distance around the inner circle once and the outer circle once...

OpenStudy (anonymous):

but doesnt the 2 already count the 8+8?

OpenStudy (whpalmer4):

No, the formula for the circumference of a circle is \(C = 2\pi r\).

OpenStudy (mathstudent55):

the 2 is from 2(pi)r, the formula for the circumference of a circle

OpenStudy (mathstudent55):

C = 2(pi)8 is only onle lap around the outer circle.

OpenStudy (anonymous):

ooooo ok